OpenStudy (anonymous):

Find the zero of the function f(x)=4x-3

OpenStudy (zzr0ck3r):

4x-3 = 0 4x=3 x = 3/4 make sense?

OpenStudy (anonymous):

why'd you drop the zero?

OpenStudy (zzr0ck3r):

4x-3 = 0 add three to both sides 4x+3-3 = 0+3 4x=3 make sense?

OpenStudy (zzr0ck3r):

@kieraoverton ?

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (anonymous):

yes. thank you

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

but I also have to graph so I need to?

OpenStudy (zzr0ck3r):

ok

OpenStudy (zzr0ck3r):

do you know what y = mx+b is ? what does m stand for and y?

OpenStudy (anonymous):

m is the slope & y is the y intercept?

OpenStudy (anonymous):

b is the y intercept

OpenStudy (anonymous):

thank you @completeidiot

OpenStudy (zzr0ck3r):

so start at the y intercept and then 4 up and 1 to the right and then start at the y intercept and go down 4 and to the left 1, because 4 = 4/1 = -4/-1 and slope is rise over run